# VecTrans: LLM Transformation Framework for Better Auto-vectorization on High-performance CPU

### v1

@@ 3.5. Optimized Case
(b) Optimized code
Figure 5. An example from Skia
4. Evaluation
4.1. Experimental Methodology
System Configuration. We perform most experiments on a Linux-based platform equipped with a 128-core Kunpeng-920 AArch64 CPU(8MB L1d cache, 8MB L1i cache, 64MB L2 cache, 128 MB L3 cache). For validation, we employ BiSheng Compiler 3.2.0.1(bisheng, ) with optimization flags -O3 -ffast-math, -Rpass=loop-vectorize, -Rpass-analysis=loop-vectorize to enable vectorization diagnostics and feedback.
Benchmarks. We evaluate a wide range of functions from TSVC_2(maleki2011evaluation, ) and real application function from Skia(google2014skia, ). The TSVC_2(Test Suite for Vectorization Compilers) benchmark suite systematically evaluates compiler capabilities in automatic vectorization through 149 test functions representing typical computational patterns. Experimental results reveal that 50 functions (approximately 33.6%) resist effective vectorization by BiSheng Compiler(bisheng, ) and mainstream compilers, primarily due to complex data dependencies, irregular memory access patterns, or unstructured control flows. The classification of these functions is shown in Table 1. These selected cases serve as critical evaluation targets, exposing limitations in current vectorization techniques while providing a standardized framework for analyzing compiler optimization boundaries and guiding algorithmic improvements.
Metrics. We consider the effectiveness of the framework from three dimensions: coverage, acceleration ratio, and iteration rounds. All of the selected functions cannot be automatically vectorized by the compiler. Through the transformation of the framework, some functions can be automatically vectorized. The proportion of these functions, that is, coverage, is our primary consideration. The speedup of the test functions that can be successfully vectorized is another considered metric reflecting the benefits of the framework. Iteration rounds reflect the resources consumed by conducting this framework for optimization, including the number of iteration rounds needed for functions that fail to be automatically vectorized, and reflect the cost of deploying the framework in the compiler.
Parameter Setup. The LLM selected for the experiment is Deepseek-V3(liu2024deepseek, ). The maximum number of iteration rounds is set to 20. When the unit test generates and analyzes whether the function is successfully vectorized (feedback), temperature is set to 1 and temperature is set to 0 in other phases. A higher temperature corresponds to increased output randomness, with the minimum temperature is zero. Set the maximum number of tokens to 4096 and retain the default values for other parameters.
Table 1. The classification of functions that cannot be vectorized in TSVC_2.
Category Function name
Could not identify reduction variable (33) s116, s123, s126, s141, s161, s211, s212, s1213, s221, s222, s231, s232, s233, s2233, s235, s2251, s256, s258, s261, s275, s277, s291, s292, s2111, s318, s3110, s3112, s321, s322, s323, s341, s342, s343
Cannot identify array bounds (2) s123, s1161
Could not determine number of loop iterations (5) s161, s277, s332, s481, s482
Instruction cannot be vectorized (2) s31111, s451
Loop contains a switch statement (1) s442
Figure 6. VecTrans generates optimized code using BiSheng Compiler(bisheng, ) on the Neon instruction set with an average speedup of 2.02x, compared with an average speedup of 1.94x for LLM-Vectorizer. The bar marked with a slash indicates that vectorization fails. It can be observed that LLM-Vectorizer and VecTrans can cover different test cases because LLM-Vectorizer is better at processing functions with conditional branches. VecTrans is better at dealing with functions with dependencies between loops.
4.2. Coverage
Our model can successfully vectorize 23 of the 50 TSCV functions, as well as a complex function from skia, coverage reached 46.2%. The coverage of LLM-Vectorizer on Neon instruction sets is only 28.8%. The functions that can be successfully vectorized are shown in Figure 6. According to the different categories shown in Table 1, among which the functions that cannot be vectorized due to unsafe dependent memory operations are most easily transformed by this framework so that they can be vectorized automatically, 10 out of 13 functions can be vectorized successfully. The functions that cannot be vectorized because they could not identify the reduction variable are more difficult to transform. Only 13 of the 33 functions can be vectorized successfully. The remaining four categories are difficult to handle. Only one test case s1161 can be vectorized successfully. Other test cases cannot generate optimization functions that can be vectorized automatically.
Table 2 shows the manner in which the LLM transformation function enables vectorization. Through these transformations, the compiler can more easily identify whether there is a dependency between loops, so that the compiler can vectorize a function that could not be vectorized originally. One of the most effective is the loop splitting technology, which can effectively eliminate many loop dependencies.
Table 2. Code Optimization Techniques and Descriptions
Optimization Description
Temporal Variable Introduce short-lived variables to store recurrent computation results, reducing dependency caused by array access..
Branch Elimination Remove conditional statements like switch or if.
In addition to the test cases that can be successfully vectorized, there are 27 test cases that cannot be successfully transformed by the framework. The conversion fails due to the following reasons:
  • The number of iteration rounds is exceeded. The upper limit of the number of iterations is set to 20. When the number of iterations exceeds 20, if the LLM still considers that the function at this time cannot be automatically vectorized, the iteration is terminated and the source code is output.
  • The LLM considers that the code has been optimized. In some cases, the LLM will ignore the vectorization information provided by the compiler. As a result, the LLM mistakenly considers that the current function can be automatically vectorized and chooses to end the iteration and output the source code.
  • The LLM considers that there is no optimization space. The LLM considers that the existing loop dependency makes the function unable to complete automatic vectorization through the compiler, and chooses to end the iteration and output the source code.
  • Generate wrong code. Among the 51 test functions, there is only one function s481, and the optimization function generated by the function s481 can pass the unit test and formal verification. However, the semantics of the optimization function is different from the original semantics, mainly because the input coverage of the unit test is limited. In addition, the formal verification tool is incomplete. This situation will output an error code, which is unacceptable, and future work will attempt to deal with it.
The functions that cannot be successfully converted by the framework have the following features: two-dimensional arrays, conditional branches, and function calls. In addition, if the loop dependency cannot be resolved, the loop cannot be vectorized successfully. The main factor is the existence of conditional branches. There are 11 test cases that cannot be transformed by the framework because of the existence of conditional branches. Second, there are 8 functions that cannot be resolved because of loop dependencies.And there are six functions that cannot be vectorized due to the difficulty of memory access analysis caused by arrays. The remaining two test cases are due to function calls, function inlining is required before the next step is carried out, which is difficult to implement at the source code level.
4.3. Speedup and Iteration Rounds
Under the Neon instruction set, a total of 24 functions are successfully vectorized using the BiSheng Compiler(bisheng, ). In comparison to the O3 optimization level, which achieves an average speedup of 2.02x and a maximum speedup of 5.21x (for function s293), two functions exhibit negative optimization outcomes and four functions show minimal or no performance improvement. This degradation in performance can be attributed to the additional overhead introduced by the transformation process, such as the creation of temporary arrays, cache misses, and other inefficiencies, which ultimately outweigh the benefits of vectorization.
The optimization framework typically necessitates a mean of 8.76 iterations per function. Functions exceeding this threshold undergo optimization cycles up to a maximum of 20 iterations, with 9 functions reaching this upper bound. Successful vectorization requires significantly fewer iterations, averaging 4.96. Each iteration round comprises two distinct Large Language Model (LLM) interactions: the feedback phase, where performance data is collected and analyzed, and the refinement phase, where the optimization function is updated based on the feedback.
Table 3. The coverage and rounds under different configurations. The Base Model means that the framework without formal verification, compiler feedback and unit test.
Configuration Coverage Iteration rounds
DeepSeek-V3/VecTrans $\mathbf{46.2\%}$ 8.760
DeepSeek-V3/LLM-Vectorizer(10.1145/3696443.3708929, ) 28.8% 13.39
DeepSeek-V3/Base Model 17.3% 5.412
Qwen2.5-32B/VecTrans 34.6% 13.940
Qwen2.5-72B/VecTrans 38.5% 12.260
DeepSeek-V3/Without Formal Verification 32.7% 8.489
DeepSeek-V3/Without Compiler Feedback 21.2% 5.213
DeepSeek-V3/Without Unit Test 25.0% 9.656
Figure 7. The graph shows the speedup of functions that can be successfully vectorized when the framework uses different LLMs. Columns with slashes indicate that vectorization cannot be successful. Three LLMs, Qwen2.5-Coder-32B-Instruct(hui2024qwen2, ), Qwen2.5-72B-Instruct(yang2024qwen2, ) and Deepseek-V3, are used here. The speedup is of the final generated optimized function compared to the original function using BishengCompiler at -O3 -ffast-math flag.
4.4. Breakdown Analysis
Table 3 shows the coverage and number of iteration rounds under different configurations. It can be seen that when a component is removed, the coverage decreases, and when all three components are removed (Base Model), the coverage reaches the lowest value. This reflects the positive effect of all three components in helping the framework complete code conversion. Compiler feedback has the greatest impact, and formal validation has the least impact.
Compiler feedback helps the entire framework in three ways as follows.
  • Helps LLM determine whether a compilation error occurs. If a compilation error occurs, the optimized function needs to be regenerated.
  • Helps LLM determine whether vectorization is successful. If vectorization has not been performed, LLM needs analyze whether vectorization opportunities exist and provide conversion suggestions.
  • Helps with LLM analysis which transformation can facilitate automatic vectorization by the compiler. The compiler gives the loop position that cannot be vectorized and the reason why it cannot be vectorized.
The removal of different components has an impact on the coverage rate and iteration rounds, but has little impact on the acceleration ratio because the code optimization methods used by the same LLM are similar and the generated code is basically the same. Except that a few functions may reduce the acceleration ratio due to redundant operations, the acceleration ratio remains the same in other cases.
Figure 8. Average speedup of optimized function generated by VecTrans under different configurations
4.5. Sensitivity Studies
The coverage rate and iteration rounds of the three different base models are shown in Table 3. It can be seen that with the decrease of model parameters, the coverage rate decreases and the number of iteration rounds increases. This is because the decrease of model parameters leads to the weakening of model capabilities, so some test functions are difficult to successfully transform. At the same time, more iteration rounds are required, and some test functions cannot output correct results within 20 iteration rounds.
Figure 7 shows the speedup of different model generation functions. The speedup of Qwen2.5-Coder-32B-Instruct is 2.9 times, and that of Qwen2.5-72B-Instinct is 2.61 times. The speedup of Deepseek-V3 is 2.54 times. Even for different LLMs, the generated functions are similar and the speedup is similar. Therefore, this framework may help identify potential compilation optimization points, that is, summarize the characteristics of the generated functions. The optimization direction of automatically vectorized can be found.
Sensitivity of different compilers. Here we explore the impact of different compilers on the optimized functions, and we test the speedup of GCC 14.2.0, Clang 20.0.0, and BiSheng Compiler 3.2.0.1(bisheng, ), respectively. The GCC and Clang versions used here are the latest versions, which represent the best compilers available today. BiSheng Compiler is a compiler optimized for Kunpeng series chips. On Kunpeng 920B, the Neon vector instruction set is used, and the speedup of functions compiled using the flag -O3 -ffast-math on different compilers is shown in Figure 8.
The average speedup of the optimized function is 1.95 times on Clang, 2.07 times on GCC, and 2.54 times on BiSheng Compiler(bisheng, ). The code generated for the s112 function is significantly worse on GCC and Clang than it is on BiSheng Compiler(bisheng, ) because the latest versions of GCC and Clang can vectorize the original s112 function. In s1161 and s242, the speedup of GCC is lower than that of the other two compilers. In s293, the speedup of Clang is lower than that of the other two compilers. This indicates that different compilers optimize functions differently. However, in most cases, the optimization of these functions is effective for all three compilers, and can achieve good results.
As shown in Figure 8, the average speedup is 2.80 times on the AVX instruction set, 2.02 times on the Neon instruction set, 2.54 times on the SVE instruction set, and on the AVX instruction set, the speedup of the s1161 function can reach 16.8 times.
It can be observed that, on different vector instruction sets, there is no obvious regularity in the speedup of the function, but the effect is better on the AVX instruction set and the SVE instruction set. The effect comes from the length of the vectorized register on the one hand, and the vectorized implementation mode on the other hand.
4.6. Summary of Results
In summary, the framework effectively transforms functions so that functions can be vectorized automatically by the compiler. In our experiments, this framework successfully vectorized 24 of the 51 functions, achieving an average speedup of 2.54 times. At the same time, this framework can improve the code generation capability of LLMs, even small models can guarantee the function generation effect. Through unit test and formal verification, the correctness of the generated code is ensured, so that the framework can be deployed in the compiler. Finally, we verify the effectiveness of the generated optimized code on different compilers and vector instruction sets.
5. Related Work
This section provides an overview of related work on LLM for compiler.
Automatic vectorization. The research on automatic vectorization has been a hot topic. Franchetti and Püschel(franchetti2008generating, ) use the term rewriting technique to vectorize small matrix kernels. Thomas et al.(thomas2024automatic, ) use equality saturation to automatically generate high-quality vectorizing compiler for digital signal processors, creating a new idea for automatic vectorization.
LLM for compiler. In recent years, there has been a lot of work using LLMs to optimize compilers, focusing on different aspects of code work. Armengol-Estapé et al.(armengol2024slade, ) demonstrate an innovative decompilation using a small language model to ensure correctness while improving code readability. Program repair is another study that has been highlighted. Bouzenia et al.(bouzenia2024repairagent, ) present a method of using LLMs to automatically repair programs, using iterative interaction, and shows how large models interact with tools .Iterative interaction framework(bi2024iterative, )(madaan2023self, ) is widely used in LLM applications to improve the ability of generation. In addition to optimizing code directly with LLM, another idea is to generate transform pass that can optimize code. Cummins et al.(cummins2024don, ) use the way of chain of thought to generate transform pass, thus improving the accuracy of LLM optimization. In addition, the LLM is used to predict the pass list, which is another way to avoid code correctness verification(cummins2023large, ). By adjusting the pass list, the performance of the program can be greatly changed.
ML for vectorization. Machine learning techniques are now being used to improve automatic vectorization. Kevin et al.(stock2012using, ) replace the cost model of automatic vectorization with a machine learning model to help the compiler make vectorization decisions. Ameer et al.(haj2020neurovectorizer, ) deploy a reinforcement learning method in the compiler to predict the optimal vectorization factor, which improves the performance of loop vectorization. Taneja et al.(10.1145/3696443.3708929, ) propose an innovative approach to vectorization, using LLM to generate source code that uses Intrinsic instructions, thus completing vectorization at the source level and verifying the semantic correctness of the vectorized code.
6. Conclusions
We implement a code transformation (source to souce) framework based on LLMs (large language models), which can effectively transform source code into a form that is more friendly for automatic vectorization. This framework gives full play to the ability of automatic vectorization of compiler. Furthermore, it uses the unit test and formal verification tools to ensure the correctness of generated code. This combination of the LLM and compiler reduces the responsibility of LLM, thus decreasing the probability of error. The proposed framework makes full use of the inference ability of LLMs, while preserving the accuracy of the traditional compilation vectorization. This combination will provide hints for future research. In future work, we will further ensure that the generated code is correct and extend the LLM to more compilation optimizations.

### v2

@@ IV-G2 Real World Application
(b) Optimized version enabling vectorization.
Figure 9: Optimization example in Skia: Eliminating dependencies via loop splitting and temporary buffering.
V Related Works
LLM for compilers. Recent work demonstrates the growing role of LLMs in compiler-related tasks. Armengol-Estapé et al.[25] demonstrate an innovative decompilation using a small language model to ensure correctness while improving code readability. Program repair is another study that has been highlighted. Bouzenia et al.[26] present a method of using LLMs to automatically repair programs, using iterative interaction, and show how large models interact with tools. Iterative interaction framework[27][7] is widely used in LLM applications to improve the ability of generation. In addition to optimizing code directly with LLM, another idea is to generate transform pass that can optimize code. Cummins et al.[28] use the way of chain of thought to generate transform pass, thus improving the accuracy of LLM optimization. In addition, the LLM is used to predict the pass list, which is another way to avoid code correctness verification[29]. By adjusting the pass list, the performance of the program can be greatly changed.
ML-based automatic vectorization. Kevin et al.[30] replace the cost model of automatic vectorization with a machine learning model to help the compiler make vectorization decisions. Ameer et al.[31] deploy a reinforcement learning method in the compiler to predict the optimal vectorization factor, which improves the performance of loop vectorization. Taneja et al.[16] propose an innovative approach to vectorization, using LLM to generate source code that uses Intrinsic instructions, thus completing vectorization at the source level and verifying the semantic correctness of the vectorized code.
Classical vectorization method. Franchetti and Püschel[32] use the term rewriting technique to vectorize small matrix kernels. Thomas et al.[33] use equality saturation to automatically generate high-quality vectorizing compiler for digital signal processors, creating a new idea for automatic vectorization.
VI Summary
We present a source code transformation framework leveraging large language models to restructure code into forms more amenable to automatic vectorization. The framework maximizes the compiler’s vectorization capabilities while employing unit tests and formal verification to ensure the correctness of the transformed code. By integrating LLMs with traditional compilers, the framework constrains the role of LLMs to inference-driven transformations, reducing the risk of semantic errors. This design preserves the reliability of conventional compiler vectorization while enhancing its applicability through learned code restructuring. We believe this approach opens new directions for integrating LLMs into compiler optimization pipelines. In future work, we plan to extend this framework to support additional optimizations and further strengthen its correctness guarantees.